OpenStudy (anonymous):

solve x^2-4x+3=0 using quadratic formula

OpenStudy (anonymous):

identify your a, b, and c in your equation: ax^2 + bx + c = 0 then plug it into the QF

OpenStudy (anonymous):

what is the quadratic equation??

jhonyy9 (jhonyy9):

do not need jut factoriz it - check two numbers what assumed will add you -4 and theyr product equal 3

jhonyy9 (jhonyy9):

just factoriz it

jhonyy9 (jhonyy9):

can you get these ?

OpenStudy (anonymous):

im lost....

jhonyy9 (jhonyy9):

hence is more easy

jhonyy9 (jhonyy9):

but with quadratic formula so do you know the discriminant

OpenStudy (anonymous):

well the quadratic formula is -b +/- *sqrt (b^2 - 4(a)(c) ------------------------ 2a So we plug it in as said before, -(-4) +/- *sqrt ((-4)^2 - 4(1)(3) ------------------------ 2(1) 4 +/- *sqrt (16 - 12) ------------------- 2 4 +/- *sqrt (4) -------------- 2 4 +/- 2 ------- 2 4 + 2 6 ------ = -- = 3 2 2 4 - 2 2 ------ = -- = 1 2 2 so the answers are x = 1 and x = 3; if anything you can enter them in the original problem to check your answer

OpenStudy (anonymous):

I attached a picture of the quadratic equation. In your problem, \[x^2-4x+3=0\] a = 1, b = -4, c = 3. So just plug in those values, and you should be able to solve your problem.